# thanks: These authors contributed equally to this workthanks: These authors contributed equally to this work

### v1

@@ II.1 The Totally Asymmetric Simple Exclusion Process
where we use the subscript 00 to indicate the system with zero paused particles or dynamical defects, the meaning of which will be clear in the following sections. In open boundary conditions, the entrance and exit boundaries fix the steady-state density which is given by the following relations:
@@ II.2 The pausing TASEP (pTASEP)
We investigate an exclusion process in which individual particles enter a paused state at a rate $k_{+}$, then exit this condition at rate $k_{-}$. As in the standard TASEP, the movement of particles occurs through hopping to the next unoccupied site, at rate $\varepsilon$. This variant of the TASEP has been introduced in [2] and then analytically investigated in [9], where the authors developed a mean-field approximation. Our ultimate goal is to study the impact of antibiotics on ribosome elongation [27]. To this end we use the pTASEP, where pausing mimics the binding of antibiotics during translation elongation. The pausing/unpausing rates $k_{+}$ and $k_{-}$ thus reflect the rates at which antibiotics bind to the ribosome, and are hence expected to be related to the intracellular antibiotic concentration $[C]$ and its binding affinity $k_{on}$ as $k_{+}=k_{on}[C]$. Once in the paused state, particles can become active again at a rate $k_{-}$, simulating antibiotics unbinding from the ribosome. Figure 2(a) provides a visual summary of these dynamics. Unless otherwise stated, and without loss of generality, we fix the particle hopping rate to $\varepsilon=1$ s-1. This is equivalent to expressing the other rates in units of $\varepsilon$.
Figure 2: Sketch of the TASEP with paused particles (pTASEP) in (a) and of the TASEP with dynamical defects (ddTASEP) in (b). In the pTASEP, particles enter and leave a paused state with the rates $k_{+}$ and $k_{-}$, respectively. In the ddTASEP (b), defects bind and unbind to holes, impeding the traffic flow on the lattice. These systems are equivalent under a particle-hole mapping ($\rho\rightarrow 1-\rho$).
@@ III.1 Mean-field approach
$f_{p}=\frac{k_{+}}{k_{+}+k_{-}}\,,\qquad f_{a}=\frac{k_{-}}{k_{+}+k_{-}}\,.$ (3)
@@
$f_{\text{J}}:=\frac{k_{-}}{k_{-}+\hat{k}_{+}}\,,\qquad\hat{k}_{+}:=k_{+}+\varepsilon\rho f_{p}\,.$
This reflects the fact that a particle contributing to the current will now longer do so after becoming paused (rate $k_{+}$) or after moving (rate $\varepsilon$) to encounter a paused particle (probability $\rho f_{p}$). In the latter case the particle, although still active, becomes blocked from advancing further.
@@
In this formulation, $f_{\text{J}}$ can be interpreted as an effective probability of a particle contributing to the current, obtained from the typical time a particle spends in the paused state $(1/k_{-})$ and from the effective time $(1/\hat{k}_{+})$ spent in its current-contributing state.
Figure 3: Comparison between stochastic simulations and mean-field approximation in (a) periodic boundary conditions and (b) open boundary conditions. When $k_{-}$ and $k_{+}$ are approximately equal to $\varepsilon$, shown in orange squares in (a) and green triangles for low density and blue diamonds for high density in (b), the agreement between simulations and theory is good. Deviations start to appear when the timescales of hopping and pausing kinetics differ significantly, as shown in dark red circles in (a) and green squares and blue circles in (b). In both plots $k_{\pm}$ are expressed in units of $\varepsilon$ and $L=1000$.
Figure 3(a) compares the prediction of this mean-field approach and the outcome of stochastic simulations (see Appendix A) with periodic boundary conditions, for two different choices of the parameters. Equation (4) clearly captures the overall behaviour seen in stochastic simulations. However, while remaining indicative of the numerical results, the analytical solution becomes less accurate when $k_{-}$ and $k_{+}$ decrease. In Section IV we will investigate more thoroughly the accuracy of mean-field predictions for a broad range of parameters.
@@ III.2 Expanding the mean-field approach to open boundary conditions
where 00 and ${L+1}$ refers to two (fictitious) reservoir sites at either end of the segment by which the lattice is extended. Their densities are chosen in order to emulate the desired entry and exit rates, respectively, assuming furthermore that the dynamics between the reservoirs and the lattice follows the same microscopical rules as in the bulk.
@@
[…]
In Fig. 3(b) we show the fundamental diagram $J(\rho)$ for the open boundary conditions, where the density $\rho$ is varied by tuning $\alpha$ and $\beta$. The left branch (LD) is obtained by varying $\alpha$ at a fixed $\beta$ and, conversely, the right branch (HD) by fixing $\alpha$ and changing $\beta$. The two branches meet in the maximal current phase, at $\rho=\rho_{max}$, i.e. where the current is limited by the bulk of the system and no longer by the boundaries. The transition from low density to maximal current occurs when $\rho_{0}=\rho_{max}$, leading to $\alpha_{crit}=f_{\text{J}}\varepsilon\rho_{max}$. Similarly, the transition between high density and maximum current takes place at $\rho_{L+1}=\rho_{max}$ which leads to $\beta_{crit}=\varepsilon(1-\rho_{max})f_{\text{J}}/f_{a}$. In Fig. 4 (a)-(b), we compare $J(\alpha)$ and $J(\beta)$ obtained following this method, for different values of $k_{-}$ and $k_{+}$. Similar to the periodic boundary condition case, the analytical predictions become worse for small values of $k_{-}/\varepsilon$ and $k_{+}/\varepsilon$. Figure 4(c) displays the phase diagram for $k_{-}/\varepsilon=k_{+}/\varepsilon=1$, the heatmap representing the outcome of simulations. Although, the fundamental diagram in Fig. 3(b), depicted as green squares and blue diamonds, show a re-scaled TASEP, the phase diagram in Fig 4(c) is shifted. This is because the pausing and unpausing dynamics act differently on the initiation and termination site which can also be seen in the aforementioned equations for $\alpha$ and $\beta$.
Figure 4: Critical values and phase diagrams. Panels (a) and (b) show the current as a function of $\alpha$ and $\beta$ respectively for $L=1000$. Similar to Fig. 3, the pTASEP captures the behavior, when the elongation rate is of the order of the pausing and unpausing rates, here represented by the dashed line. On the contrary, the prediction, depicted by the solid line, becomes worse for lower values of $k_{-}$ and $k_{+}$ (expressed in units of $\varepsilon$) compared to the elongation rate. Panel (c) depicts the entire phase diagram for $k_{-}=k_{+}=\varepsilon=1$. Straight lines correspond to the predictions of $\alpha_{crit}$ and $\beta_{crit}$ while the heatmap represents the outcome of the simulated current.
@@ IV.1 The transcription and translation parameter space
Transcription. Single-molecule experiments on transcription in vitro have revealed that RNAPs enter and exit a paused state with rates that are between 0.07 - 0.15 s-1 and approximately 1 s-1, respectively [2]. The elongation rate of RNAPs ranges from 20 to 80 nucleotides per second, resulting in a timescale separation between pausing and unpausing dynamics of one to three orders of magnitude [29, 30, 31]. Moreover, the ddTASEP has been used in the literature to describe the impact of DNA blockages on RNAP traffic. These dynamical blockages can be any regulatory proteins binding and unbinding the DNA template, ranging from histones [32] to methylated-DNA binding protein [33]. For instance, the MeCP2 binding protein has a $k_{\pm}~{}0.04/s$, to be compared with the RNAP elongation rate mentioned above, i.e. $\varepsilon\sim 20-80$ nucleotides/s.
@@
In the translation process, the elongation rate is between 10-20 amino acids per second [35]. The binding and unbinding times of the aforementioned antibiotics are several orders of magnitude lower then the elongation rate, thus leading to a timescale separation between the translation and antibiotics dynamics. For instance, chloramphenicol has a binding rate constant of ($k_{on}=5.6\times 10^{-4}\mu M^{-1}s^{-1}$) and an unbinding rate of ($k_{-}=1.23\times 10^{-3}s^{-1}$[17]. As typical sublethal doses of antibiotics range from $2$ $\mu M$ to $12$ $\mu M$ [36], the timescales are up to five orders of magnitude apart.
Thus, in order to investigate transcription and translation in a biologically relevant regime, we need to study the regime $k_{-},k_{+}\ll\varepsilon$.
@@ IV.2 Periodic boundary conditions
Mimicking the aforementioned effect of antibiotics, we have simulated the pTASEP in periodic boundary conditions by varying $k_{+}$ over several orders of magnitude.
@@
By comparing the model to the outcome of simulations we notice that the mean-field prediction fails, not only quantitatively but also qualitatively, when decreasing $k_{+}$, see Fig. 5(a)-(c). We observe that the discrepancy between the mean-field prediction given by Eq. (11) and simulations arises when there is likely no paused particle on any of the finite number of sites. The expected number of paused particles is given by $N_{p}:=f_{p}\rho L$, which depends on the rates $k_{\pm}$, on the particle density and the lattice length. The impact of the blockages imposed by paused particles, and thus Eq. (11), is relevant only when $N_{p}$ is at least comparable to $1$, which is when the total number of particles is sufficiently large ($\rho L\geq f_{p}^{-1}$). Otherwise, we expect finite sizes to be relevant.
@@
In Fig. 5 we show $J(\rho)$ by varying the microscopic pausing rate $k_{+}$, panels (a)-(c), or by changing the lattice size, panels (d)-(f). We notice that, when the number of expected paused particles $N_{p}$ is order $1$ or smaller the current exceeds the pTASEP mean-field predictions (compare the simulation points and the orange curve). For instance, at the peaks of panels (b) and (e), $N_{p}\sim 1$. With paused particles being scarce, Fig. 5(a)-(b)-(d)-(e), the system behaves as a standard TASEP (i.e. without paused particles) for extensive amounts of time, with a current therefore exceeding the mean-field predictions. For larger densities, the probability of observing paused particles increases ($N_{p}>1$) and the mean-field argument captures the outcome of the simulations.
@@
Figure 5: Relevance of the finite size effects in the $J(\rho)$ relation. The pTASEP and single cluster solution are shown in blue and orange respectively, whereas the numerical simulation is in grey. The figures (a) to (c) show the emergence of finite size effects by varying $k_{+}$ from $10^{-6}$, $10^{-4}$ and $10^{-2}$, respectively, while keeping $L=250$ and $k_{-}=10^{-3}$. The figures (d) to (f), show the same effect emerging by changing the system size $L$, while keeping $k_{-}=10^{-3}$ and $k_{+}=10^{-4}$ fixed.
@@
Understanding the role of the system size is therefore crucial: finite-size effects in the pTASEP are intrinsically related to the pausing-unpausing dynamics via $N_{p}$, and are significant in biologically relevant regimes. To investigate finite-size effects, we juxtapose Fig. 5(d)-(f) by using the same dynamical parameters as in the central panel (b), but with progressively larger lattice sizes (up to $L=1000$). We observe that the ‘bump’ in the current disappears for large lattices, whereas it is amplified in small lattices. Additionally, deviations from the mean-ield predictions become more pronounced in smaller lattices. Thus, simulations clearly show that there is a size dependent behavior, and that mean-field approaches greatly underestimate the current when finite size is relevant (i.e. when $N_{p}\leq 1$). We address this situation by developing a two-term equation, with one term describing the current produced in the absence of pauses, weighted by the probability $P_{0}$ of having no paused particles on the lattice, and the other describing the current produced by the system in the complementary case, weighted by $1-P_{0}$. Since the particle number $\rho L$ is constant, as well as the average ratio of active particles $f_{a}$ (at steady state), $P_{0}$ is given by
$P_{0}=(f_{a})^{\rho L}=\left(\frac{k_{-}}{k_{-}+k_{+}}\right)^{\rho L}$ (10)
@@
$J=P_{0}J_{0}+(1-P_{0})J_{c}\,,$ (11)
where the current is $J_{0}$ when no particles are in the paused state, and $J_{\text{MF}}$ when at least one particle is paused. We will see that, in the regime of interest of rare and long pauses, the transition of a particle to the paused state implies the formation of large clusters.
In the following, we will first develop a single-cluster approximation and obtain a solution for $J_{c}$ and $J_{0}$ to better understand the regime dominated by finite size effects, and compare this to results from the literature. Eventually we provide a more phenomenological solution extending our approach to all regimes.
IV.2.1 Single-cluster approximation for $J_{\text{MF}}$ and $J_{0}$
When the particle movement is much faster than the pausing dynamics (i.e. in the biologically relevant regime $k_{-},k_{+}\ll\varepsilon$) it is reasonable to first assume that all particles cluster together behind a particle in the paused state. In this state, unpausing events of the leading particle determine an intermittent current $J_{c}$ established by the detaching fronts. When finite-size effects are relevant ($N_{p}$ is order $1$), there is a non-negligible probability of dissolving the complete cluster, which effectively goes to zero for high density. In this case, the system will relax from the ‘cluster state’ to the standard TASEP state (with no particles paused), with a particle current $J_{0}$. In the single-cluster steady state, paused particles are separated by $d$ active particles as shown in Fig. 6(a). When the leading particle becomes unpaused, $d+1$ particles detach from the cluster, representing an effective density $(d+1)/L$. Due to the aforementioned timescale separation between the two dynamics and the modest lattice lengths of the systems of our interest, we make the hypothesis that these particles travel the $L(1-\rho)$ empty sites without becoming paused and rejoin the cluster from the other end. This sub-cluster detachment occurs intermittently, with a typical time between events given by $1/k_{-}$, the inverse of the unpausing rate of the particle at the front of the cluster. Putting all this together, in the clustered regime the mean particle current reads
$J_{c}=k_{-}(1-\rho)(d+1)\,.$ (12)

### v2

@@ II.1 The Totally Asymmetric Simple Exclusion Process
where we use the subscript 0 to indicate the system with zero paused particles or dynamical defects, the meaning of which will be clear in the following sections. In open boundary conditions, the entrance and exit boundaries fix the steady-state density which is given by the following relations:
@@ II.2 The pausing TASEP (pTASEP)
We investigate an exclusion process in which individual particles enter a paused state at a rate $k_{p}$, then exit this condition at rate $k_{u}$. As in the standard TASEP, the movement of particles occurs through hopping to the next unoccupied site, at rate $\varepsilon$. This variant of the TASEP has been introduced in [2] and then analytically investigated in [9], where the authors developed a mean-field approximation. Our ultimate goal is to study the impact of antibiotics on ribosome elongation [27]. To this end we use the pTASEP, where pausing mimics the binding of antibiotics during translation elongation. The pausing/unpausing rates $k_{p}$ and $k_{u}$ thus reflect the rates at which antibiotics bind to the ribosome, and are hence expected to be related to the intracellular antibiotic concentration $[C]$ and its binding affinity $k_{on}$ as $k_{p}=k_{on}[C]$. Once in the paused state, particles can become active again at a rate $k_{u}$, simulating antibiotics unbinding from the ribosome. Figure 2(a) provides a visual summary of these dynamics. Unless otherwise stated, and without loss of generality, we fix the particle hopping rate to $\varepsilon=1$ s-1. This is equivalent to expressing the other rates in units of $\varepsilon$.
Figure 2: Sketch of the TASEP with paused particles (pTASEP) in (a) and of the TASEP with dynamical defects (ddTASEP) in (b). In the pTASEP, particles enter and leave a paused state with the rates $k_{p}$ and $k_{u}$, respectively. In the ddTASEP (b), defects bind and unbind to holes, impeding the traffic flow on the lattice. These systems are equivalent under a particle-hole mapping ($\rho\rightarrow 1-\rho$).
@@ III.1 Mean-field approach
$f_{p}=\frac{k_{p}}{k_{p}+k_{u}}\,,\qquad f_{a}=\frac{k_{u}}{k_{p}+k_{u}}\,.$ (3)
@@
$f_{\text{J}}:=\frac{k_{u}}{k_{u}+\hat{k}_{p}}\,,\qquad\hat{k}_{p}:=k_{p}+\varepsilon\rho f_{p}\,.$
This reflects the fact that a particle contributing to the current will now longer do so after becoming paused (rate $k_{p}$) or after moving (rate $\varepsilon$) to encounter a paused particle (probability $\rho f_{p}$). In the latter case the particle, although still active, becomes blocked from advancing further.
@@
In this formulation, $f_{\text{J}}$ can be interpreted as an effective probability of a particle contributing to the current, obtained from the typical time a particle spends in the paused state $(1/k_{u})$ and from the effective time $(1/\hat{k}_{p})$ spent in its current-contributing state.
Figure 3: Comparison between stochastic simulations and mean-field approximation in (a) periodic boundary conditions and (b) open boundary conditions. When $k_{u}$ and $k_{p}$ are approximately equal to $\varepsilon$, shown in orange squares in (a) and green triangles for low density and blue diamonds for high density in (b), the agreement between simulations and theory is good. Deviations start to appear when the timescales of hopping and pausing kinetics differ significantly, as shown in dark red circles in (a) and green squares and blue circles in (b). In both plots $k_{p}$ and $k_{u}$ are expressed in units of $\varepsilon$ and $L=1000$.
Figure 3(a) compares the prediction of this mean-field approach and the outcome of stochastic simulations (see Appendix A) with periodic boundary conditions, for two different choices of the parameters. Equation (4) clearly captures the overall behaviour seen in stochastic simulations. However, while remaining indicative of the numerical results, the analytical solution becomes less accurate when $k_{u}$ and $k_{p}$ decrease. In Section IV we will investigate more thoroughly the accuracy of mean-field predictions for a broad range of parameters.
@@ III.2 Expanding the mean-field approach to open boundary conditions
where 0 and ${L+1}$ refers to two (fictitious) reservoir sites at either end of the segment by which the lattice is extended. Their densities are chosen in order to emulate the desired entry and exit rates, respectively, assuming furthermore that the dynamics between the reservoirs and the lattice follows the same microscopical rules as in the bulk.
@@
$\rho_{max}:=-\chi+\sqrt{\chi^{2}+\chi}\hskip 11.38109pt\text{and}\hskip 11.38109pt\chi:=\frac{k_{p}+k_{u}}{\varepsilon f_{p}}.$ (9)
In Fig. 3(b) we show the fundamental diagram $J(\rho)$ for the open boundary conditions, where the density $\rho$ is varied by tuning $\alpha$ and $\beta$. The left branch (LD) is obtained by varying $\alpha$ at a fixed $\beta$ and, conversely, the right branch (HD) by fixing $\alpha$ and changing $\beta$. The two branches meet in the maximal current phase, at $\rho=\rho_{max}$, i.e. where the current is limited by the bulk of the system and no longer by the boundaries. The transition from low density to maximal current occurs when $\rho_{0}=\rho_{max}$, leading to $\alpha_{crit}=f_{\text{J}}\varepsilon\rho_{max}$. Similarly, the transition between high density and maximum current takes place at $\rho_{L+1}=\rho_{max}$ which leads to $\beta_{crit}=\varepsilon(1-\rho_{max})f_{\text{J}}/f_{a}$. In Fig. 4 (a)-(b), we compare $J(\alpha)$ and $J(\beta)$ obtained following this method, for different values of $k_{u}$ and $k_{p}$. Similar to the periodic boundary condition case, the analytical predictions become worse for small values of $k_{u}/\varepsilon$ and $k_{p}/\varepsilon$. Figure 4(c) displays the phase diagram for $k_{u}/\varepsilon=k_{p}/\varepsilon=1$, the heatmap representing the outcome of simulations. Although, the fundamental diagram in Fig. 3(b), depicted as green squares and blue diamonds, show a re-scaled TASEP, the phase diagram in Fig 4(c) is shifted. This is because the pausing and unpausing dynamics act differently on the initiation and termination site which can also be seen in the aforementioned equations for $\alpha$ and $\beta$.
Figure 4: Critical values and phase diagrams. Panels (a) and (b) show the current as a function of $\alpha$ and $\beta$ respectively for $L=1000$. Similar to Fig. 3, the pTASEP captures the behavior, when the elongation rate is of the order of the pausing and unpausing rates, here represented by the dashed line. On the contrary, the prediction, depicted by the solid line, becomes worse for lower values of $k_{u}$ and $k_{p}$ (expressed in units of $\varepsilon$) compared to the elongation rate. Panel (c) depicts the entire phase diagram for $k_{u}=k_{p}=\varepsilon=1$. Straight lines correspond to the predictions of $\alpha_{crit}$ and $\beta_{crit}$ while the heatmap represents the outcome of the simulated current.
@@ IV.1 The transcription and translation parameter space
Transcription. Single-molecule experiments on transcription in vitro have revealed that RNAPs enter and exit a paused state with rates that are between 0.07 - 0.15 s-1 and approximately 1 s-1, respectively [2]. The elongation rate of RNAPs ranges from 20 to 80 nucleotides per second, resulting in a timescale separation between pausing and unpausing dynamics of one to three orders of magnitude [29, 30, 31]. Moreover, the ddTASEP has been used in the literature to describe the impact of DNA blockages on RNAP traffic. These dynamical blockages can be any regulatory proteins binding and unbinding the DNA template, ranging from histones [32] to methylated-DNA binding protein [33]. For instance, the MeCP2 binding protein has a $k_{p}$ and $k_{u}\approx 0.04/s$, to be compared with the RNAP elongation rate mentioned above, i.e. $\varepsilon\sim 20-80$ nucleotides/s.
@@
In the translation process, the elongation rate is between 10-20 amino acids per second [35]. The binding and unbinding times of the aforementioned antibiotics are several orders of magnitude lower then the elongation rate, thus leading to a timescale separation between the translation and antibiotics dynamics. For instance, chloramphenicol has a binding rate constant of ($k_{on}=5.6\times 10^{-4}\mu M^{-1}s^{-1}$) and an unbinding rate of ($k_{u}=1.23\times 10^{-3}s^{-1}$[17]. As typical sublethal doses of antibiotics range from 222 $\mu M$ to 121212 $\mu M$ [36], the timescales are up to five orders of magnitude apart.
Thus, in order to investigate transcription and translation in a biologically relevant regime, we need to study the regime $k_{u},k_{p}\ll\varepsilon$.
@@ IV.2 Periodic boundary conditions
Mimicking the aforementioned effect of antibiotics, we have simulated the pTASEP in periodic boundary conditions by varying $k_{p}$ over several orders of magnitude.
@@
By comparing the model to the outcome of simulations we notice that the mean-field prediction fails, not only quantitatively but also qualitatively, when decreasing $k_{p}$, see Fig. 5(a)-(c). We observe that the discrepancy between the mean-field prediction given by Eq. (11) and simulations arises when there is likely no paused particle on any of the finite number of sites. The expected number of paused particles is given by $N_{p}:=f_{p}\rho L$, which depends on the rates $k_{p}$ and $k_{u}$, on the particle density and the lattice length. The impact of the blockages imposed by paused particles, and thus Eq. (11), is relevant only when $N_{p}$ is at least comparable to 111, which is when the total number of particles is sufficiently large ($\rho L\geq f_{p}^{-1}$). Otherwise, we expect finite sizes to be relevant.
@@
In Fig. 5 we show $J(\rho)$ by varying the microscopic pausing rate $k_{p}$, panels (a)-(c), or by changing the lattice size, panels (d)-(f). We notice that, when the number of expected paused particles $N_{p}$ is order 111 or smaller the current exceeds the pTASEP mean-field predictions (compare the simulation points and the orange curve). For instance, at the peaks of panels (b) and (e), $N_{p}\sim 1$. With paused particles being scarce, Fig. 5(a)-(b)-(d)-(e), the system behaves as a standard TASEP (i.e. without paused particles) for extensive amounts of time, with a current therefore exceeding the mean-field predictions. For larger densities, the probability of observing paused particles increases ($N_{p}>1$) and the mean-field argument captures the outcome of the simulations.
@@
Figure 5: Relevance of the finite size effects in the $J(\rho)$ relation. The pTASEP and single cluster solution are shown in blue and orange respectively, whereas the numerical simulation is in grey. The figures (a) to (c) show the emergence of finite size effects by varying $k_{p}$ from $10^{-6}$, $10^{-4}$ and $10^{-2}$, respectively, while keeping $L=250$ and $k_{u}=10^{-3}$. The figures (d) to (f), show the same effect emerging by changing the system size $L$, while keeping $k_{u}=10^{-3}$ and $k_{p}=10^{-4}$ fixed.
@@
Understanding the role of the system size is therefore crucial: finite-size effects in the pTASEP are intrinsically related to the pausing-unpausing dynamics via $N_{p}$, and are significant in biologically relevant regimes. To investigate finite-size effects, we juxtapose Fig. 5(d)-(f) by using the same dynamical parameters as in the central panel (b), but with progressively larger lattice sizes (up to $L=1000$). We observe that the ‘bump’ in the current disappears for large lattices, whereas it is amplified in small lattices. Additionally, deviations from the mean-ield predictions become more pronounced in smaller lattices. Thus, simulations clearly show that there is a size dependent behavior, and that mean-field approaches greatly underestimate the current when finite size is relevant (i.e. when $N_{p}\leq 1$). We address this situation by formulating a two-term expression: one term accounts for the current produced in the absence of paused particles (the unpaused state), weighted by the probability $P_{0}$, and the other term captures the current in the presence of at least one paused particle (the paused state), weighted by the complementary probability $1-P_{0}$. Since the particle number $\rho L$ is constant, as well as the average ratio of active particles $f_{a}$ (at steady state), $P_{0}$ is given by
$P_{0}=(f_{a})^{\rho L}=\left(\frac{k_{u}}{k_{u}+k_{p}}\right)^{\rho L}$ (10)
@@
$J=P_{0}J_{0}+(1-P_{0})J_{\text{p}}\,,$ (11)
where $J_{0}$ denotes the current in the unpaused state (no paused particle on the lattice), and $J_{\text{p}}$ corresponds to the current when at least one particle is in the paused state. As we will see, in the regime of rare and long pauses, a single pausing event typically leads to the formation of large clusters.
In the following, we will first develop a single-cluster approximation and obtain a solution for $J_{p}$ and $J_{0}$ to better understand the regime dominated by finite size effects, and compare this to results from the literature. Eventually we provide a more phenomenological solution extending our approach to all regimes.
IV.2.1 Single-cluster approximation for $J_{\text{p}}$ and $J_{0}$
[…]
$J_{p}=k_{u}(1-\rho)(d+1)\,.$ (12)